Margaret is 14 years old. two years ago, she was twice as old as her brother Tony.
a) how old is Tony now?
b) when will they both be the same age?

Answer by ankor@dixie-net.com(22740)

Margaret is 14 years old. two years ago, she was twice as old as her brother Tony.
:
let b = brothers present age
14 - 2 = 2(b-2)
12 = 2b - 4
12 + 4 = 2b
16 = 2b
a) How old is Tony now
b = 8 yrs old now
:
b) when will they both be the same age? Never, a dumb question